OpenStudy (revircs):

how would i differentiate 3^(cos(xsinx))?

OpenStudy (solomonzelman):

Use the chain rule twice. First, tell me what is the derivative of 3^x ?

OpenStudy (revircs):

That would be ln(3) x 3^x

OpenStudy (solomonzelman):

Yes, \(\large\color{black}{ \displaystyle \frac{ d }{dx} 3^x=3^x\ln(3) }\)

OpenStudy (solomonzelman):

Then, by the Chain Rule principal you get: \(\large\color{black}{ \displaystyle \frac{ d }{dx} 3^{f(x)}=3^{f(x)}\ln(3)\times f'(x) }\)

OpenStudy (solomonzelman):

So then: \(\Large\color{black}{ \displaystyle \frac{ d }{dx} 3^{\cos\left(\sin x\right)}=3^{\cos\left(\sin x\right)}\ln(3)\times ? }\)

OpenStudy (solomonzelman):

You just need to find that chain rule derivative for \(\Large\color{black}{ \displaystyle {\cos\left(\sin x\right)} }\)

OpenStudy (revircs):

what about with an x in front of the sin, like \[3^{\cos(xsin(x))}\]

OpenStudy (solomonzelman):

Oh, there is an x in front of sin x? Then, \(\large\color{black}{ \displaystyle \frac{ d }{dx} 3^{\cos\left(x\sin x\right)}=3^{\cos\left(x\sin x\right)}\ln(3)\times \left\{\frac{d}{dx}~\cos\left(x\sin x\right)\right\} }\)

OpenStudy (revircs):

would the derivative of cos(xsinx) be -sin(xsinx)*x*cos(x)?

OpenStudy (solomonzelman):

x*sinx will go by the product rule

OpenStudy (solomonzelman):

what is the derivative of x•sin(x) ?

OpenStudy (revircs):

1*cos(x), I believe.

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle \frac{ d}{dx}(f\cdot g)=f'g+fg' }\)

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle \frac{ d}{dx}(x\cdot \sin x)=(x')\cdot \sin x+x \cdot (\sin x') }\)

OpenStudy (revircs):

oh, so just sin(x)+(x*cos(x))?

OpenStudy (solomonzelman):

yes, \(\large\color{black}{ \displaystyle \frac{ d}{dx}(x\cdot \sin x)= \sin x+x \cos x}\)

OpenStudy (solomonzelman):

And therefore: \(\large\color{black}{ \displaystyle \frac{ d}{dx}~\sin(x\cdot \sin x)=\cos(x\cdot \sin x)\times (\sin x+x \cos x)}\)

OpenStudy (solomonzelman):

If you ever don't get something, please ask...

OpenStudy (solomonzelman):

And lastly, therefore: \(\normalsize\color{black}{\displaystyle \frac{ d}{dx}~3^{\Large\sin(x\cdot \sin x)}=\ln(3)\cdot 3^{\Large\sin(x\cdot \sin x)}\cdot \color{red}{\cos(x\cdot \sin x)\times (\sin x+x \cos x)}}\)

OpenStudy (revircs):

Thank you so much!

OpenStudy (solomonzelman):

yw